OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4

OpenStudy (agreene):

-2(x+3)=-2(x+1)-4 Start by distributing -2 -2x-6=-2x-2-4 Combine like terms -2x-6=-2x-6 This is an obviously true statement. if you wish, you can extend this by adding 2x and 6 to both sides getting: 0=0

OpenStudy (precal):

Solution is all real numbers.